Break-even analysis for a service company Rotelco is one of the largest digital wireless service providers in the United States. In a recent year, it had approximately 100 direct subscribers (accounts) that generated revenue of $43,800. Costs and expenses for the year were as follows: Cost of revenue $19,700 Selling, general, and administrative expenses 13,600 Depreciation 4,800 Assume that 80% of the cost of revenue and 20% of the selling, general, and administrative expenses are variable to the number of direct subscribers (accounts). In part (a) and (b), round all interim calculations and final answers to one decimal place.

a. What is Rotelco's break-even number of accounts, using the data and assumptions above? Round to the nearest whole number. accounts

b. How much revenue per account would be sufficient for Rotelco to break even if the number of accounts remained constant? Round to the nearest dollar. $ per account
Answer #1
Calculation of fixed costs, variable costs and variable cost per unit:
Cost Cost of Revenue Selling, general, and
administrative expenses
Depreciation
Expense (a)                       19,700                               13,600                  4,800
Variable cost (%) (b) 80% 20% 0%
Variable cost ($) ( c) = (a)*(b)                       15,760                                 2,720                         -  
Fixed Cost (d) = (a)-(c )                          3,940                               10,880                  4,800
Number of subscribers (e)                             100                                     100                         -  
Variable cost per unit ($) (f) = ( c)/(e)                       157.60                                 27.20                         -  
Part a
Calculation of Break even number of accounts
Break even (number of accounts) = Total Fixed costs / Contribution margin per unit
Sales revenue (a)                       43,800
Number of subscribers (b)                             100
Sales revenue per unit ( c) = (a)/(b)                       438.00
Less : Variable costs per unit
Cost of Revenue (d)                      -157.60
Selling, general, and
administrative expenses
(e)                        -27.20
Contribution margin per unit (f) = ( c)-(d)-(e)                       253.20
Fixed costs
Cost of Revenue                          3,940
Selling, general, and
administrative expenses
                      10,880
Depreciation                          4,800
Total Fixed costs                       19,620
Break even (number of accounts) = Total Fixed costs / Contribution margin per unit
= 19620/253.2
= 77.5
Hence 78 subscribers are needed to break even
Part b
Calculation of Break even number of accounts
Break even (revenue) = Total Fixed costs / Profit Volume Ratio
Profit Volume Ratio = (Contribution margin/sales revenue)*100
Sales revenue (a)                       43,800
Number of subscribers (b)                             100
Sales revenue per unit ( c) = (a)/(b)                       438.00
Less : Variable costs per unit
Cost of Revenue (d)                      -157.60
Selling, general, and
administrative expenses
(e)                        -27.20
Contribution margin per unit (f) = ( c)-(d)-(e)                       253.20
Profit volume ratio (g) = [(f) / ( c)]*100 57.8%
Fixed costs
Cost of Revenue                          3,940
Selling, general, and
administrative expenses
                      10,880
                                                                              -                            4,800
Total Fixed costs                       19,620
Break even (revenue) = Total Fixed costs / Profit Volume Ratio
= 19620 / 57.8%
= $33,945
Revenue per account (based on actual number of subscribers) = 33945 / 100
= $339.5
Revenue per account (based on break even number of subscribers) = 33945 / 78
= $435.20